OpenStudy (anonymous):

A change machine changes dollar bills into quarters and nickels. If you receive 12 coins after inserting a dollar bill, how many nickels did you receive?

OpenStudy (anonymous):

it should be 11

OpenStudy (anonymous):

well its not so

OpenStudy (anonymous):

2 quarters, 10 nickels. Give me a minute to show the math.

OpenStudy (anonymous):

25x+5y=100 x+y=12 x=12-y Substitute back in after solving for x. 25(12-y)+5y=100 300-25y+5y=100 300-100=20y y=10 Back into the original problem: x=12-10 x=2